Lillie is three years older than her sister Shieley. Lillie's brother is six years younger than their sister Shirley. The sum of all their ages is 42.
Please help with writing this equation

Answer by mathdad1965(8)

Let Shirley = x

Let Lillie = x + 3

Lillie's brother = x - 6

Sum = 42

x + (x + 3) + (x - 6) = 42